Double T resort has two departments, which are rooms and F&B. Rooms and F&B departments’ revenues are $280,000 and $130,000 respectively. Also, the VCs are $200,000 and $90,000 for rooms and F&B department respectively. Suppose that room’s revenue has increased by 14.00% and there is a 12.00% decrease in VC for both departments. If the FC is $160,000, how much reduction in total revenue at the breakeven point has Double T experienced before and after the changes in both revenues and VCs?  

Group of answer choices

$199,023 reduction

$176,179 reduction

$237,561 reduction

$203,822 reduction
Answer #1

Contribution = Sales - Variable cost

Room Contribution = 280000 - 200000 = 80000

F&B contribution = 130000 - 90000 = 40000

Overall contribution = 80000 + 40000 = 120000

Overall sales = 280000 + 130000 = 410000

Overall contribution margin = 120000 / 410000 = 29.2683%

Break even sales = Fixed costs / Overall contribution margin = 160000 / 29.2683% = 546666

After revision

Room contribution = (280000 *1.14 - 200000 * 0.88) = 143200

F&B contribution = 130000 - 90000 *0.88 = 50800

Overall sales = 280000*1.14 + 130000 = 449200

Overall contribution = 143200 + 50800 = 194000

Contribution margin = 194000 / 449200 = 43.1879%

Revised break even = 160000 / 43.1879% = 370473

Reduction in break even sales = 546666 - 370473 = 176193

Correct Answer: B

Note : difference may be due to rounding.